OpenStudy (anonymous):

(Medal + Fan) A line includes the points (1,0) and (12,9). What is the equation of the line in point slope form?

OpenStudy (anonymous):

\[y- y1 = m(x-x1) \] I know the formula

OpenStudy (anonymous):

First find slope \[m=\frac{ (y2-y1) }{ (x2-x1)}\]

OpenStudy (anonymous):

\[m=\frac{ \left( y_2 - y_1 \right) }{ \left( x_2 - x_1 \right) }\] this will get you the slope and... \[y-y_1 = m \left( x-x_1 \right)\] is the point slope form of a line

OpenStudy (anonymous):